OpenStudy (anonymous):

PLEASE HELP? A ball is launched from a slingshot. Its height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. After 1 second, the ball is 148 feet in the air, after 2 seconds the ball is 272 feet in the air. Find the height in feet of the ball after 6 seconds in the air?

OpenStudy (turingtest):

A quadratic equation has the form\[Ax^2+Bx+C=h(x)\]is there anything in the problem that indicates the initial height of the ball before launch is zero?

OpenStudy (anonymous):

I'm not sure how to use that form to get the answer. From reading the problem I do not see anything that indicates the initial height of the ball. Just what it is at after 1 second, and then after 2 seconds.

OpenStudy (logan13):

y is this q yellow

iYuko (iyuko):

Qualified helper. He used OwlBucks

OpenStudy (logan13):

oh

OpenStudy (igreen):

Isn't it \(\sf H(t) = -16t^2 + vt + s\)?

OpenStudy (turingtest):

I guess we can say that h(x) represents the change in height from its initial position. If we can do that we can say the equation is of the form\[Ax^2+Bx=h(x)\]we know that when x=1, h(x)=148, so we can get an equation from that.

OpenStudy (anonymous):

I'm confused :/

OpenStudy (turingtest):

when x=2, h(x)=272, so plugging that in gives another equation we can then solve the system of equations to find A and B

OpenStudy (turingtest):

@iGreen I think this question assumes no knowledge of physics

OpenStudy (igreen):

Oh, okay.

OpenStudy (turingtest):

@paigeally are you expected to know the acceleration of gravity? Is it given anywhere?

OpenStudy (anonymous):

No it doesn't expect physics, or gravity. It's just a quadratic function question where I have to find a function that'll work to find out what 6 seconds is. It's from Alegbra 2

OpenStudy (turingtest):

Then I will say that h(x) represents the change in height, and so a quadratic formula that expresses this is\[h(x)=Ax^2+Bx\]When x=1, h(x)=148, so we can write\[h(1)=A(1)^2+B(1)^2=148\]

OpenStudy (anonymous):

I searched for help for the same question, and a similar problem with different words and number came up, and it showed a way to figure it out but I couldn't understand it.

OpenStudy (anonymous):

How can I use that formula you gave me to figure the answer? I'm not sure what goes in A,B, or C

OpenStudy (turingtest):

but do you understand what I said in my last post? this gives us an equation with two unknowns (A and B)

OpenStudy (turingtest):

We have to assume C=0, or else the problem cannot be solved. For that reason I don't like the wording of the question.

OpenStudy (michele_laino):

I think that the height can be modeled by the subsequent function: \[\Large h\left( x \right) = a{x^2} + b\]

OpenStudy (anonymous):

Can I post the similar solution to a similar question, and maybe that will help you explain what I need to do?

OpenStudy (turingtest):

sure

OpenStudy (anonymous):

A ball is launched from a sling shot. After 1 second, the ball is 121 feet in the air. After 2 seconds, it's 224 feet in the air. Using quadratic functions, find the height, in feet, of the ball after 3 seconds in the air. 0a +0b +c = 0 ; c = 0 ---------------------------- a + b = 121 ; a = 121-b ----------------------------- 4a +2b = 224 4(121-b) + 2b = 224 484 -4b +2b = 224 484 - 224 = 2b 260/2 = b ; b = 130 ------------------------------ a + 130 = 121 a = 121 - 130 ; a = -9 ------------------------------------ our quadratic formula becomes: y = -9x^2 +130x y = -9(3)^2+130(3) y = -81 + 390 y = 309 ; the height is 309 ft after 3 seconds.

OpenStudy (turingtest):

ok good, that is the method I am doing

OpenStudy (turingtest):

they also assumed c=0 so again, we have the formula\[Ax^2+Bx=h(x)\]we know that when x=1, h(x)=148 (in other words, h(1)=148) so plug those numbers into this equation and you get \[A(1)^2+B(1)=148\]which simplifies to \[A+B=148\]so far so good?

OpenStudy (anonymous):

Yes so far I follow

OpenStudy (turingtest):

ok, now we also know that h(2)=272, so again we can plug x=2 into our equation, and we get\[A(2)^2+B(2)=272\]which simplifies to what?

OpenStudy (anonymous):

A+b=272 ?

OpenStudy (turingtest):

\(A*2^2=4A\)

OpenStudy (turingtest):

\(B*2=2B\)

OpenStudy (anonymous):

Alright I see how you got that

OpenStudy (turingtest):

good, so we have two equations:\[A+B=148\\4A+2B=272\]this is a system of two equations with two unknowns, which means it can be solved. Do you know how to do this using substitution?

OpenStudy (anonymous):

No, I don't. Math isn't my forte.

OpenStudy (turingtest):

well, we want to get it so one of the equations has only one variable in it. An easy way to do that in this case is to solve for either A or B in the first equation, then substitute that into the second equation. Can you solve the first equation for B?

OpenStudy (anonymous):

It could be any number though, how do I know what the right one is?

OpenStudy (turingtest):

I mean solve \(A+B=148\) for B. In other words, get it into the form where B is alone on one side and A and 148 are on the other, i.e.\[B=...\]

OpenStudy (anonymous):

I'm not following, do you mean pick two numbers to fill in for A and B?

OpenStudy (turingtest):

no, I mean subtract A from both sides

OpenStudy (turingtest):